# DAFE: LLM-Based Evaluation Through Dynamic Arbitration for Free-Form Question-Answering

### v1

@@ 4.1 Alignment with human evaluation
LLMs Tasks EM BS DeepSeek Llama GPT Mixtral Mistral DAFE
Llama AmbigQA 0.518 0.283 0.897 0.888 0.844 0.824 0.858 0.911
HotpotQA 0.577 0.498 0.885 0.877 0.899 0.820 0.832 0.953
NQ-Open 0.381 0.437 0.797 0.833 0.793 0.816 0.738 0.927
TriviaQA 0.281 0.564 0.460 0.547 0.439 0.396 0.299 0.684
GPT AmbigQA 0.561 0.252 0.951 0.944 0.897 0.861 0.853 0.967
HotpotQA 0.604 0.300 0.807 0.953 0.973 0.873 0.933 0.987
NQ-Open 0.453 0.218 0.809 0.884 0.824 0.824 0.829 0.956
TriviaQA 0.335 0.364 0.594 0.650 0.401 0.580 0.467 0.775
Mixtral AmbigQA 0.546 0.337 0.896 0.896 0.781 0.909 0.887 0.951
HotpotQA 0.546 0.349 0.920 0.940 0.933 0.859 0.940 0.973
NQ-Open 0.371 0.301 0.825 0.879 0.728 0.899 0.815 0.913
TriviaQA 0.317 0.390 0.661 0.625 0.605 0.678 0.436 0.764
Mistral AmbigQA 0.599 0.254 0.893 0.893 0.893 0.893 0.860 0.953
HotpotQA 0.605 0.383 0.903 0.937 0.902 0.895 0.937 0.958
NQ-Open 0.484 0.291 0.797 0.851 0.838 0.878 0.840 0.953
TriviaQA 0.467 0.239 0.754 0.758 0.725 0.645 0.470 0.854
As evidenced by consistently high Macro F1 scores in Table 2, DAFE maintains a strong alignment with human judgment. This represents a substantial improvement over individual model performance, where individual judges generally revealed varying levels of agreement with human evaluation. LLM-as-a-judge approach generally works better with larger more powerful models. This is particularly noticeable in DeepSeek and GPT which achieve higher Macro-F1 scores (0.97-0.98) across AmbigQA, HotpotQA, and NQ-Open compared to smaller models. This reveals an important scaling law in evaluation capability (Kaplan et al., 2020; Zheng et al., 2024; OpenAI et al., 2024). However, we also found that the most advanced models are not always guaranteed to be the best evaluators. We observed slightly comparable performance through small open-source Mistral-7B. For instance, when evaluating candidate Mixtral-8x7B on AmbigQA, Mistral-7B as-a-judge outperformed (0.944) judge GPT-3.5-turbo (0.891). Regardless, we observe relatively lower Macro-F1 scores for all LLM judges in TriviaQA.
Interestingly, despite EM’s deviation from the human majority (see Figure 2 and Table 6), lexical matching EM typically accomplishes better alignment with human evaluation on instance-level in Table 2 than neural-based BERTScore. EM’s strict and conservative nature leads to lower overall performance, but its high-precision characteristics ensure that when it identifies a match, it strongly aligns with human judgment. In contrast, BERTScore takes a more lenient approach to semantic matching. Although this leniency produces higher raw scores, it introduces more false positives, consequently reducing instance-level agreement with human judgments. This pattern emerges clearly in many models and tasks such as when evaluating Llama-3.1-70B on AmbigQA, EM shows a raw score of 42.3% but achieves a Macro-F1 of 0.744, while BERTScore indicates a higher raw score of 63.0% but a lower Macro-F1 of 0.641.
Table 2: Macro-F1 scores of various evaluators applied to different candidate LLMs and associated tasks. Higher scores indicate better performance. DAFE consistently achieves the highest Macro-F1 across all evaluated settings.
Evaluators
LLMs Tasks EM BS DeepSeek Llama GPT Mixtral Mistral DAFE
Llama AmbigQA 0.744 0.641 0.948 0.944 0.922 0.912 0.929 0.955
HotpotQA 0.778 0.745 0.942 0.939 0.949 0.910 0.916 0.976
NQ-Open 0.653 0.718 0.898 0.916 0.896 0.907 0.869 0.964
TriviaQA 0.612 0.782 0.726 0.772 0.717 0.695 0.640 0.842
GPT AmbigQA 0.792 0.622 0.976 0.972 0.949 0.930 0.927 0.984
HotpotQA 0.794 0.623 0.903 0.977 0.987 0.936 0.966 0.993
NQ-Open 0.703 0.606 0.904 0.942 0.911 0.911 0.914 0.978
TriviaQA 0.646 0.681 0.796 0.824 0.700 0.789 0.730 0.887
Mixtral AmbigQA 0.760 0.666 0.948 0.948 0.891 0.955 0.944 0.975
HotpotQA 0.761 0.657 0.960 0.970 0.966 0.930 0.970 0.987
NQ-Open 0.650 0.649 0.912 0.939 0.863 0.950 0.908 0.956
TriviaQA 0.625 0.695 0.829 0.812 0.803 0.838 0.716 0.882
Mistral AmbigQA 0.792 0.622 0.947 0.947 0.947 0.947 0.930 0.977
HotpotQA 0.796 0.673 0.951 0.969 0.951 0.947 0.969 0.979
NQ-Open 0.726 0.639 0.898 0.925 0.919 0.939 0.920 0.976
TriviaQA 0.718 0.608 0.925 0.879 0.863 0.822 0.735 0.927
Figure 3: Heatmap illustrating the performance of four individual LLM judges on HotpotQA. Each cell value represents percentages (%). Rows represent predicted outcomes (P: Positive, N: Negative), while columns represent actual outcomes. See Appendix C for full results.
4.2 Analysis
In our experiments, candidate LLMs generated 7,500 outputs for the given tasks, with each evaluator producing 7,500 corresponding evaluations. We randomly sampled 100 error cases (50 false positives and 50 false negatives) from each evaluator to understand their behavior. Given EM had only 11 false positives, we included all of them in our analysis. Due to space constraints, we moved the detailed analysis of EM and BERTScore to Appendix C and focused exclusively on the LLM-as-a-judge method here.
LLM-based evaluators demonstrate strong abilities in recognizing semantic variations while maintaining the core meaning, especially when assessing responses that use different terminology or structural approaches to convey the same information. For instance, in the evaluation examples, evaluators correctly identified that “Salma Hayek” and “Salma Hayek Pinault” refer to the same individual, acknowledging the semantic equivalence despite differences in phrasing. Similarly, when assessing responses that use different terms for the same entity, such as recognizing “Nick Fury, Agent of S.H.I.E.L.D.” as part of the broader “Marvel” universe, the evaluators effectively maintain the core meaning and contextual relevance. Their explanations show systematic assessment patterns that combine multiple evaluation criteria including factual accuracy, logical coherence, and contextual relevance.
LLMs are prone to hallucination in justification (Zhang et al., 2023), where they fabricate reasoning to support their evaluations, produce detailed but incorrect explanations, or reference non-existent criteria or standards. In LLM judges, false positives and negatives (e.g., see Figure 3) often result from overlooking critical distinctions between candidate LLM outputs and failing to account for the specificity required by the reference answer. This pattern is particularly noticeable in Mistral 7B, where the model disregards the ground truth and provides evaluations influenced by unknown factors. For example, when evaluating candidate GPT-3.5’s response “The foreign minister of Germany who signed the Treaty of Versailles was Hermann Müller.” which is correct according to the reference answer “Hermann Müller” and human evaluation, Mistral 7B as-a-judge incorrectly marked this response as false and fabricated reasoning “Hermann Müller was the Chancellor of Germany, not the Foreign Minister. The Foreign Minister of Germany who signed the Treaty of Versailles was Gustav Stresemann.” in support of its decision. The same problem can also be attributed to inconsistent evaluations. Because when Mistral 7B acted as a candidate for the same question, its response to the question is completely different: “The Treaty of Versailles was signed by Matthias Erzberger, a German politician who served as the President of the German National Assembly at the time”. There are also alternative interpretations of this issue, such as ambiguity in the question, but we leave a deeper exploration of these aspects to future work.
We observe a different pattern in some judges, specifically, GPT-3.5 and Mixtral 8x7B which focuses more on specificity. This approach shifts the evaluation towards false negatives by missing semantically similar but structurally different answers. We found many cases when such evaluators failed to account for valid variations in phrasing or granularity, focusing instead on rigid adherence to the reference answer. Compounding these issues are reasoning errors within the evaluators’ own explanations, which often contain fabrications, circular logic, or overconfident assertions. By insisting on correctness derived strictly from the reference, evaluators disregard valid alternative perspectives and can even mischaracterize or invert the facts in their attempts to justify their decisions. This dynamic leaves little room for nuance or ambiguity, and it pushes the evaluation process away from fair, context-sensitive assessment toward rigid, and sometimes inaccurate, verdicts.
Verbosity (Ye et al., 2024) emerges as a subtle source of bias, where more elaborate answers are sometimes overrated simply due to their detail and fluency, while concise yet correct responses are undervalued. This misplaced emphasis leads to irrelevant judgment criteria, such as praising the presence of irrelevant information or penalizing perfectly valid but succinct answers. We also found that LLM-based judges encounter challenges in multiple reference answers and more open-ended questions. This confusion is especially pronounced in the TriviaQA where the diversity and flexibility of valid responses present challenges for the judges’ ability to consistently recognize and evaluate a range of correct answers.
We found several temporal limitations in LLM-based evaluators. Although most of our datasets are older and the evaluator models are relatively up-to-date, we still observed instances where references to recent events, newly emerging terminology, or evolving contexts were misinterpreted. The FreshQA dataset (Vu et al., 2023), being recent, serves as a valuable testbed for assessing these temporal deficiencies. As shown in Table 3, LLM-based evaluators indicate deviation from human judgment on FreshQA compared to tasks that rely on older information, such as HotpotQA. Specifically, in dynamic or time-sensitive contexts, we found that LLM judges tend to hallucinate by consistently classifying candidate model responses as True, even when incorrect. For example, when presented with the question: “On what date did the Patriots last play the Miami Dolphins?” the LLM-generated response states: ‘‘The last time the New England Patriots played the Miami Dolphins was on January 1, 2023, during the NFL regular season.” Despite the correct reference answer being “November 24, 2024” the LLM evaluator not only failed to recognize the inaccuracy but also hallucinated an erroneous justification, stating: “The proposed answer correctly states the date the New England Patriots last played the Miami Dolphins as January 1, 2023, which matches the information provided.”
Table 3: Performance (in Macro F1) of LLM judges on FreshQA.
Evaluators
LLMs DeepSeek Llama GPT Mixtral Mistral DAFE
DeepSeek 0.714 0.692 0.715 0.614 0.724 0.830
Llama 0.801 0.835 0.737 0.817 0.730 0.917
GPT 0.659 0.695 0.824 0.780 0.746 0.891
Mixtral 0.732 0.708 0.779 0.738 0.703 0.936
Mistral 0.687 0.665 0.802 0.818 0.723 0.880
Figure 4: Disagreement rates between primary judges (Llama+Mistral) across candidate LLMs and tasks.
4.3 Disagreements between primary judges
Figure 4 shows that disagreements between our primary judges, Llama-3.1 70B and Mistral 7B, mainly occur in the TriviaQA and FreshQA, with disagreement rates reaching 20.3% and 44.3%, respectively. Interestingly, higher disagreement rates between primary judges create a greater opportunity for DAFE to refine evaluations. As depicted in Figure 4, FreshQA (31.3% for Llama-70B, 39.7% for Mistral-7B) demonstrates the highest disagreement, allowing DAFE to improve Macro F1 scores (see Table 3).
4.4 Impact of arbitration
Our proposed arbitration approach significantly enhanced evaluation performance by resolving disputes through an independent judge, GPT-3.5-turbo (see Figure 5). Notably, in the AmbigQA, Macro F1 scores advanced from 72.9% to 86.6%, and Cohen’s Kappa increased from 0.467 to 0.773 (see Figure 7). These improvements highlight the pivotal role of the arbitrator in ensuring reliable and consistent evaluation outcomes.
Figure 5: Comparison of Macro F1 scores before and after arbitration (see Appendix C for more results).
@@ 5 Related work
Evaluation of natural language generation has traditionally relied on supervised signal-based metrics such as EM which evaluates the exact lexical match between generated outputs and reference answers. Despite its simplicity and efficiency, EM overlooks semantically equivalent variations, often penalizing accurate responses that use different phrasing (Wang et al., 2024a; Kamalloo et al., 2023). Other commonly used metrics including BLEU Papineni et al. (2002) and ROUGE (Lin, 2004) primarily focus on n-gram overlap with human written reference texts. Despite their widespread use, these metrics have significant limitations in capturing semantic subtleties and contextual relevance (Zhang et al., 2020). To address the limitations of conventional metrics, various model-based methods such as BERTScore Zhang et al. (2020) offer semantically informed evaluation. However, even BERTScore and similar embedding-based methods struggle to effectively evaluate open-ended generation Zheng et al. (2024); Sun et al. (2022).
Recent advances in LLMs have unlocked new opportunities for automatic and context-aware evaluation (Li et al., 2024b; Chiang and Lee, 2023; Zheng et al., 2024). A key strength of LLM-based evaluators lies in their ability to operate in reference-free settings, where evaluation does not rely on pre-defined answers but instead leverages subjective criteria such as helpfulness, relevance, and coherence. This capability makes LLM evaluators particularly well-suited for assessing tasks where multiple valid responses exist or where human-like judgment is required (Li et al., 2024a). For instance, LLMs are frequently used in subjective evaluations such as pairwise comparison (“Which response is better?”) or single-response scoring (“How good is this response based on criteria X?” (Verga et al., 2024; Chan et al., 2024). LLM-based evaluators are specifically effective for tasks like summarization, where subjective criteria are central to evaluation (Liu et al., 2023). However, they are less effective for fact-based tasks such as free-form question-answering, where responses are either correct or incorrect and require explicit verification against reference answers.
Furthermore, LLM-based evaluators face several challenges, particularly in ensuring consistency and fairness (Ye et al., 2024; Khan et al., 2024). In reference-free settings, the absence of a definitive ground truth increases the risk of bias in evaluations (Ye et al., 2024; Kim et al., 2024; Huang et al., 2024a). Common biases include positional bias, where LLMs may favor responses based on their order (Zheng et al., 2024; Khan et al., 2024), verbosity bias, which favors longer or more detailed responses (Huang et al., 2024b), and self-enhancement bias, where models may disproportionately prefer their own outputs (Zheng et al., 2024). These biases can distort evaluations and undermine the reliability of the results.
@@ 6 Conclusion
We present DAFE, a framework designed to evaluate free-form question-answering by leveraging LLMs. Our findings demonstrate that individual LLM judges are reliable alternatives to traditional lexical and neural-based metrics, offering closer alignment with human evaluations. However, relying solely on individual judges poses challenges including inherent biases and prompt sensitivity, which can affect evaluation performance. DAFE addresses these challenges through a dynamic arbitration mechanism. This design achieves near-perfect agreement with human evaluations, establishing DAFE as a trustworthy and reliable framework for evaluating open-ended language generation tasks. In the future, we aim to explore DAFE by excluding reference answers and integrating LLM agents with tools-interacting capabilities for evaluation.
@@ 7 Limitations
We acknowledge certain limitations in our study. The accuracy of evaluations depends on the quality and clarity of reference answers, which serve as the basis for determining correctness. Inconsistent or ambiguous references could affect evaluation outcomes. Similarly, this study primarily uses binary verdicts which might overlook detailed aspects of responses that could be captured through more comprehensive evaluation criteria. Furthermore, while we conducted an error analysis of LLM judges and automatic metrics, there may be error cases that were not identified during our manual review, leaving gaps in understanding the full spectrum of evaluation inaccuracies. Finally, our study focuses exclusively on English, and the applicability of our approach to other languages, particularly morphologically rich or resource-scarce ones, remains unexplored.
@@ Appendix A Free-form Question-Answering
In our experiments, we include AmbigQA (Min et al., 2020), FreshQA (Vu et al., 2023), HotpotQA (Yang et al., 2018), Natural Questions (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017).
@@
Appendix B Human evaluation
This section provides detailed guidelines for human annotators responsible for evaluating the outputs of candidate LLMs. The goal is to ensure consistency and objectivity across all evaluations. These guidelines provide clear instructions for assessing each model’s response based on its alignment with the reference answer and contextual relevance.
B.1 Guidelines
Dear Evaluator,
B.2 Inter human annotator agreement
We calculate Fleiss’ Kappa ($\kappa$(Fleiss and Cohen, 1973) to assess inter-rater reliability among human annotators. Table 4 and 5 show the inter-annotator agreement across models and tasks.
@@ B.2 Inter human annotator agreement
The results demonstrate high reliability, with Fleiss’ Kappa scores consistently above 0.93 for most tasks. The highest agreement is observed in Mixtral evaluations on HotpotQA ($\kappa=0.996$), and GPT on NQ-Open ($\kappa=0.990$). In FreshQA, which shows lower Kappa scores, the agreement among annotators remains high including 99.3% in GPT and 98.0% in Mixtral.
@@
Appendix C Additional results
This section provides further results and analysis of conventional metrics and LLM-based evaluators. Table 6 illustrates the overall performance of candidate LLMs obtained through various evaluators. Unlike lexical matching and neural-based metrics, each LLM-as-a-judge indicates overall performance close to the human majority. Automatic metrics like EM severely underestimate the candidate LLMs’ performance. On the other hand, BERTScore tends to overestimate the performance.

### v2

@@ 4.2 Error analysis
Candid. Llama GPT Mistral CLEV Disagr. (%)
Llama 0.835 0.737 0.730 0.917 31.3
GPT 0.695 0.824 0.746 0.891 44.3
Mixtral 0.708 0.779 0.703 0.936 37.3
Mistral 0.665 0.802 0.723 0.880 39.7
Table 3: Macro-F1 of judges on FreshQA and corresponding disagreement rates. Higher disagreement indicates a more frequent need for a resolving judge (CLEV).
@@ 5 Related work
Recent advances in LLMs have unlocked new opportunities for automatic and context-aware evaluation (li2024llmsasjudgescomp; chianglee2023large; 10.5555/3666122.3668142). Early work primarily focused on subjective evaluation tasks, such as pairwise comparison for ranking model outputs or single-response scoring for open-ended generation (chan2024chateval; 10.5555/3666122.3668142; chen2024humansllmsjudgestudy). These approaches leverage LLMs’ strong language understanding and reasoning capabilities to assess qualities like coherence, fluency, and helpfulness. However, such subjective criteria and continuous scoring schemes are less suitable for evaluating objective QA (badshah2025taletoolaugmented), where free-form answers are expected to be factually correct with respect to a reference answer (krumdick2025no; shi2024judgingjudg).
To address this gap, recent work has explored reference-grounded LLM-as-judge approaches that instruct LLMs to directly verify candidate answers against gold references and output binary True/False verdicts (kamalloo-etal-2023-evaluating; NEURIPS2023_f323d594). However, this line of work also highlights key challenges: individual LLM judges can be inconsistent across questions, may over-rely on parametric knowledge when references are incomplete, and often require additional judges or majority voting to stabilize decisions (khan2024debating). Furthermore, single judges often lack robustness to adversarial inputs and may hallucinate justifications that appear plausible but are factually incorrect (hu2024llmbasedeval).
To mitigate the reliability issues of single-judge systems, multi-judge setups have been explored (badshah-sajjad-2025-reference; verga2024replacing). These approaches employ multiple diverse LLM judges and aggregate their decisions, typically through majority voting, to improve reliability and alignment with human evaluation. badshah-sajjad-2025-reference demonstrated that using multiple models reduces individual model biases and improves agreement with human judgments. However, these approaches typically rely on fixed ensembles, where all judges are invoked for every instance, regardless of whether the initial judges already agree. This design leads to substantial and unnecessary computational overhead, especially when disagreements are rare. Our work builds on these foundations that invoke a third judge only when needed, achieving the reliability benefits of multi-judge voting while saving compute.
@@ 6 Conclusion
We studied the challenge of evaluating free-form QA, where traditional metrics fail to capture semantic correctness and single LLM judges may be inconsistent. To address this, we introduced CLEV, a lightweight multi-judge framework that preserves the reliability of majority voting while avoiding its computational cost. By invoking a third judge only when necessary, CLEV achieves substantial to near-perfect agreement with human evaluation while reducing redundant evaluation calls by 80–95% across tasks. This makes LLM-based evaluation more scalable for large-scale benchmarking. While our study focuses on binary correctness, future work can extend CLEV to capture dimensions such as partial correctness. Looking forward, an important direction is reference-free evaluation, where judges are supported by external verifiers (e.g., retrieval) to assess correctness in settings where explicit references are lacking.
@@ 7 Limitations
We acknowledge certain limitations in our study: 1) The accuracy of evaluations depends on the quality and clarity of reference answers, which serve as the basis for determining correctness. Incorrect or ambiguous references could affect evaluation outcomes. 2) While we conducted an error analysis of LLM judges and automatic metrics, there may be error cases that were not identified during our manual review, leaving gaps in understanding the full spectrum of evaluation inaccuracies. 3) We acknowledge that our current study is limited to English QA datasets. Extending CLEV to multilingual settings is a promising and important future direction. 4) Since the evaluation is binary, which is standard for free-form QA, signals around partial correctness, fluency, reasoning, and justification are not captured. To tackle this, we also collected detailed rationales that justify each verdict. In our analysis, we leveraged these rationales to capture various insights. We acknowledge the limitation of binary verdicts in representing partial correctness, but we emphasize that this design choice reduces the subjectivity inherent in continuous scoring and promotes clearer agreement between judges. 5) CLEV assumes that primary judges maintain a minimum reliability threshold (based on $\kappa$ and Macro-F1) established during calibration. These thresholds require recalibration when adopting new judge models, domains, or languages. Thus, applying CLEV to new settings may involve additional upfront calibration cost.
Acknowledgment
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), Canada Foundation for Innovation (CFI), and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
@@ Appendix A Free-form Question-Answering
In our experiments, we include AmbigQA (minetal2020ambigqa), FreshQA (vu2023freshllms), HotpotQA (yang2018hotpotqa), Natural Questions (kwiatkowskietal2019), and TriviaQA (joshi2017triviaqa).
@@
Appendix B Baselines
Exact Match (EM):
For our selected datasets and also free-form QA tasks, EM serves as a standard lexical matching metric to evaluate candidate LLM performance (izacardgrave2021; 105555; gou2024criticllms). Due to the verbose nature of LLM-generated responses, we adapt EM to classify an answer as correct if any golden answer $r_{i}\in R$ appears within the generated response $\bar{y}$ (i.e., $r_{i}\subseteq\bar{y}$), rather than requiring complete strict string equality (i.e., $\bar{y}=r_{i}$).
BERTScore:
We use BERTScore (ZhangKWWA20) which measures similarity by comparing contextualized word embeddings derived from a pre-trained BERT model. This enables the evaluation to focus on semantic correctness rather than exact lexical matches. As BERTScore is based on continuous values between -1 and 1, we set a threshold of $\tau=0.5$ to convert continuous similarity scores into binary 0 and 1. The purpose of this conversion is to allow direct comparison with other evaluation methods. For our implementation, we use the microsoft/deberta-xlarge-mnli [^3] https://huggingface.co/microsoft/deberta-xlarge-mnli model (he2021deberta).
Majority voting
This uses three fixed LLM judges to independently evaluate each instance. The final decision is determined by a simple majority across the three verdicts. Unlike CLEV, which selectively invokes the third judge only in cases of disagreement, this method uniformly engages all judges, leading to a higher computational cost.
Appendix C Human evaluation
This section provides detailed guidelines for human annotators responsible for evaluating the outputs of candidate LLMs. The goal is to ensure consistency and objectivity across all evaluations (yu-etal-2024-latent). These guidelines provide clear instructions for assessing each model’s response based on its alignment with the reference answer and contextual relevance.
@@ Appendix C Human evaluation
C.1 Guidelines
Dear Evaluator,
Appendix D Evaluation Metrics
We compute Fleiss’ Kappa ($\kappa$(fleiss1973equivalence) and percent agreement to assess inter-rater reliability among human annotators. Similarly, we use Cohen’s kappa (mchugh2012interrater) to find the agreement between each evaluator and the human majority to obtain instance-level comparison. Due to the high-class imbalance in TriviaQA, kappa scores can be misleadingly low despite high raw agreement - a known limitation called the “kappa paradox” (cicchetti1990high). Therefore, we treat the evaluation as a binary classification task where we consider each evaluator’s predictions against the human majority and report Macro-F1 scores which give equal weight to both classes regardless of their frequency in the selected random samples.
To quantify the efficiency of our approach, we report the disagreement rate between the two primary judges that indicates how often the third model is required, thereby revealing the reduction in third-model usage compared to always employing three judges. Formally,
$\text{Disagreement rate (\%)}=\left(\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[V_{i_{1}}\neq V_{i_{2}}]\right)\times 100$
where $N$ is the total number of evaluation instances and $\mathbb{I}[\cdot]$ is the indicator function that equals 1 when the condition is satisfied and 0 otherwise.
D.1 Inter-human annotator agreement
We calculate Fleiss’ Kappa ($\kappa$(fleiss1973equivalence) and percent agreement to assess inter-rater reliability among human annotators.
Fleiss’ Kappa is defined as:
$\kappa=\frac{\bar{P}-P_{e}}{1-P_{e}},$
where $\bar{P}$ is the average observed agreement among annotators, and $P_{e}$ is the expected agreement by chance.
Percent agreement is calculated as:
$\text{Percent Agreement}=\left(\frac{\text{Agreements}}{\text{Total Annotations}}\right)\times 100$
Table 4 and 5 show the inter-annotator agreement across models and tasks. The results demonstrate high reliability, with Fleiss’ Kappa scores consistently above 0.93 for most tasks. The highest agreement is observed in Mixtral evaluations on HotpotQA ($\kappa=0.996$), and GPT on NQ-Open ($\kappa=0.990$). In FreshQA, which shows lower Kappa scores, the agreement among annotators remains high including 99.3% in GPT and 98.0% in Mixtral.
Appendix E Additional results
This section provides further results and analysis of conventional metrics and LLM-based evaluators.
As evidenced by high Cohen’s kappa and Macro F1 scores in Table 6 and 7, CLEV maintains a strong alignment with human evaluation. This represents a substantial improvement over individual model performance, where individual judges generally showed varying levels of agreement with human evaluation. Overall, LLM-as-a-judge works better with larger models. This is particularly noticeable in Llama and GPT, which achieve greater performance across AmbigQA, HotpotQA, and NQ-Open compared to smaller models. This indicates an important scaling law in evaluation capability (kaplan2020scalinglawsneurallanguage; 10.5555/3666122.3668142; openai2024gpt4technicalreport). However, we also found that the most advanced models are not always guaranteed to be the best evaluators. We observed slightly comparable performance through the small open-source Mistral7B. For instance, when evaluating candidate Mixtral 8x7B on AmbigQA (see Table  7), Mistral 7B as-a-judge outperformed (0.944) judge GPT-3.5-turbo (0.891). Regardless, we observe relatively lower Macro-F1 scores for all LLM judges in TriviaQA.
Evaluators
LLMs Tasks EM BS Llama GPT Mixtral Mistral CLEV
Llama AmbigQA 0.518 0.283 0.888 0.844 0.824 0.858 0.911
HotpotQA 0.577 0.498 0.877 0.899 0.820 0.832 0.953
NQ-Open 0.381 0.437 0.833 0.793 0.816 0.738 0.927
TriviaQA 0.281 0.564 0.547 0.439 0.396 0.299 0.684
GPT AmbigQA 0.561 0.252 0.944 0.897 0.861 0.853 0.967
HotpotQA 0.604 0.300 0.953 0.973 0.873 0.933 0.987
NQ-Open 0.453 0.218 0.884 0.824 0.824 0.829 0.956
TriviaQA 0.335 0.364 0.650 0.401 0.580 0.467 0.775
Mixtral AmbigQA 0.546 0.337 0.896 0.781 0.909 0.887 0.951
HotpotQA 0.546 0.349 0.940 0.933 0.859 0.940 0.973
NQ-Open 0.371 0.301 0.879 0.728 0.899 0.815 0.913
TriviaQA 0.317 0.390 0.625 0.605 0.678 0.436 0.764
Mistral AmbigQA 0.599 0.254 0.893 0.893 0.893 0.860 0.953
HotpotQA 0.605 0.383 0.937 0.902 0.895 0.937 0.958
NQ-Open 0.484 0.291 0.851 0.838 0.878 0.840 0.953
TriviaQA 0.467 0.239 0.758 0.725 0.645 0.470 0.854
Table 6: Cohen’s Kappa scores displaying the agreement levels of individual and multiple (CLEV) evaluators with human judgments across candidate models and tasks.
Evaluators
LLMs Tasks EM BS Llama GPT Mixtral Mistral CLEV
Llama AmbigQA 0.744 0.641 0.944 0.922 0.912 0.929 0.955
HotpotQA 0.778 0.745 0.939 0.949 0.910 0.916 0.976
NQ-Open 0.653 0.718 0.916 0.896 0.907 0.869 0.964
TriviaQA 0.612 0.782 0.772 0.717 0.695 0.640 0.842
GPT AmbigQA 0.792 0.622 0.972 0.949 0.930 0.927 0.984
HotpotQA 0.794 0.623 0.977 0.987 0.936 0.966 0.993
NQ-Open 0.703 0.606 0.942 0.911 0.911 0.914 0.978
TriviaQA 0.646 0.681 0.824 0.700 0.789 0.730 0.887
Mixtral AmbigQA 0.760 0.666 0.948 0.891 0.955 0.944 0.975
HotpotQA 0.761 0.657 0.970 0.966 0.930 0.970 0.987
NQ-Open 0.650 0.649 0.939 0.863 0.950 0.908 0.956
TriviaQA 0.625 0.695 0.812 0.803 0.838 0.716 0.882
Mistral AmbigQA 0.792 0.622 0.947 0.947 0.947 0.930 0.977
HotpotQA 0.796 0.673 0.969 0.951 0.947 0.969 0.979
NQ-Open 0.726 0.639 0.925 0.919 0.939 0.920 0.976
TriviaQA 0.718 0.608 0.879 0.863 0.822 0.735 0.927
Table 7: Macro-F1 scores of individual and multiple (CLEV) evaluators applied to different candidate LLMs and associated tasks.
Unlike lexical matching and neural-based metrics, each LLM-as-a-judge shows overall performance close to the human majority. The proposed CLEV method consistently achieves comparable or slightly better alignment with the human majority. Conventional metrics such as EM severely underestimate the candidate LLMs’ performance. Contrarily, BERTScore tends to overestimate the performance except in some cases such as when evaluating Llama on AmbigQA and NQ-Open.